suppose that for some function g, g(x+5)=3x+2. find g(-2)
Answer by Gogonati(855)

g(x+5)=g(-2) if x+5=-2 or x=-7 Since g(-7)=3x+2 we get 3(-7)+2=-19
As a conclusion g(-2)=-19.